OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4 plz help

OpenStudy (anonymous):

Not math section dude

OpenStudy (anonymous):

lol soz i thought i was

OpenStudy (nurali):

3(x + 1) = -2(x - 1) - 4 3x+3=-2x+2-4 3x+3=-2x-2 3x+2x=-2-3 5x=-5 x=-1

OpenStudy (anonymous):

lol thxs i got it right

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

NP